Assume that IQ scores are normally distributed with a mean of 100 and a standard deviation of 15.

Find the percentage of the population which has an IQ score between 104 and 112.
Answer #1

Solution :

Given that ,

mean =   = 100

standard deviation = = 15

P(104< x <112 ) = P[(104-100) / 15< (x - ) / < (112-100) / 15)]

= P( 0.27< Z <0.8 )

= P(Z < 0.8) - P(Z <0.27 )

Using z table   

= 0.7881-0.6064

= 0.1817

answer=18.17%